# Building on Efficient Foundations: Effectively Training LLMs with Structured Feedforward Layers

### v1

@@ B.5 Complete results table
Architecture Model FFN Model Training Loss PPL Size (M) Size (M) FLOPs (G) Tokens (B) Steps (K) FLOPs Transformer-s 110 56.62 262.9 2.2 4K 1.69e+18 3.2569 25.97 LowRank 73.95 20.94 189.8 2.2 4K 1.22e+18 3.3748 29.22 LowRank 3.0 6K 1.69e+18 3.2928 26.92 LowRank 2.2 5K 1.69e+18 3.2866 26.75 BlockDense 73.95 20.94 189.8 2.2 4K 1.22e+18 3.3731 29.17 BlockDense 3.0 6K 1.69e+18 3.2982 27.06 BlockDense 2.2 5K 1.69e+18 3.2856 26.73 BlockShuffle 73.95 20.94 189.8 2.2 4K 1.22e+18 3.3994 29.95 BlockShuffle 3.0 6K 1.69e+18 3.3218 27.71 BlockShuffle 2.2 5K 1.69e+18 3.3011 27.14 Transformer-m 335.08 201.33 788.7 6.7 13K 1.55e+19 2.9062 18.29 LowRank 202.43 68.68 517.0 6.7 13K 1.01e+19 3.0251 20.60 LowRank 10.2 19K 1.54e+19 2.9359 18.84 LowRank 6.7 18K 1.55e+19 2.9310 18.75 BlockDense 198.67 64.91 509.3 6.7 13K 1.00e+19 3.0371 20.85 BlockDense 10.4 20K 1.55e+19 2.9491 19.09 BlockDense 6.7 18K 1.55e+19 2.9420 18.95 BlockShuffle 202.43 68.68 517.0 6.7 13K 1.01e+19 3.0501 21.12 BlockShuffle 10.2 19K 1.54e+19 2.9627 19.35 BlockShuffle 6.7 18K 1.55e+19 2.9525 19.15 Transformer-l 729.11 452.98 1646.9 14.6 28K 7.03e+19 2.6594 14.29 LowRank 430.66 154.53 1035.6 14.6 28K 4.42e+19 2.7527 15.69 LowRank 23.3 44K 7.03e+19 2.6917 14.76 LowRank 14.6 39K 7.01e+19 2.6850 14.66 BlockDense 430.66 154.53 1035.6 14.6 28K 4.42e+19 2.7570 15.75 BlockDense 23.3 44K 7.03e+19 2.6946 14.80 BlockDense 14.6 39K 7.01e+19 2.6941 14.79 BlockShuffle 430.66 154.53 1035.6 14.6 28K 4.42e+19 2.7735 16.01 BlockShuffle 23.3 44K 7.03e+19 2.7053 14.96 BlockShuffle 14.6 39K 7.01e+19 2.7104 15.04 Transformer-xl 1274.14 805.31 2814.3 25.5 49K 2.10e+20 2.5226 12.46 LowRank 743.56 274.73 1727.7 25.5 49K 1.29e+20 2.6062 13.55 LowRank 41.5 79K 2.10e+20 2.5464 12.76 LowRank 25.5 70K 2.10e+20 2.5539 12.86 BlockDense 728.49 259.65 1696.8 25.5 49K 1.27e+20 2.6204 13.74 BlockDense 42.2 80K 2.10e+20 2.5590 12.92 BlockDense 25.5 71K 2.10e+20 2.5637 12.98 BlockShuffle 743.56 274.73 1727.7 25.5 49K 1.29e+20 2.6254 13.81 BlockShuffle 41.5 79K 2.10e+20 2.5623 12.97 BlockShuffle 25.5 70K 2.10e+20 2.5678 13.03
@@
Figure 12: Performance between dense Transformer, Structured FFN (BlockDense and BlockShuffle, LowRank in 1(b)) with 32% parameters either trained with self-guided training or more tokens across four sizes. Circle size indicates model FLOPs. To enable straightforward comparison, we controlled their training FLOPs to be the same.
@@ Appendix C Implementation details
Model We employ the basic decoder-only Transformer architecture [41, 2] but enhance it with advanced components like Rotary Embedding [42] and SentencePiece Tokenizer [43], as used in recent Llama models [51]. We use a vocabulary size of 32,000 and enable tied word embeddings. The model employs layer normalization and the GELU activation function, with biases included in the architecture. We use random weight gaussian initialization with an initializer range of 0.02.
Structured FFN We limit the replacement of efficient linear layers only to FFN modules. This is because the linear layers in the attention module serve specific purposes, which can complicate analyses, and their efficient alternatives have been well-studied in the literature [52, 8, 22].
@@
Dataset We use the RefinedWeb dataset [44], a carefully curated subset of CommonCrawl, processed for optimal filtering and deduplication. It provides 600B tokens available for public use. As it is a substantial dataset, we shuffle, extract, and tokenize it in advance. To manage memory efficiently, the token IDs are stored using np.memmap, preventing the need to simultaneously load all data into CPU memory. We set the maximum sequence length to be 1024. Obeying these, we randomly select 0.5B tokens for validation and reserve additional tokens for training. Following scaling laws [45], we allocate tokens at 20 times the number of each baseline model’s parameters.
Training We use A100 80G GPU to do the training and evaluation and adopt mixed precision technique (bfloat16 and float32) supported by torch.cuda.amp to accelerate training. For details, we employ the AdamW optimizer with 0.1 weight decay and [0.9, 0.999] betas and use the cosine annealing learning rate scheduler with 10% warm-up and $0.1$ times the minimum value. Learning rates are first swept for baseline models and applied consistently in subsequent experiments. Dropout is set to 0.0, and gradient clipping to 1.0. We measure the training FLOPs like Megatron [48], including all matrix multiplications. The configuration of each baseline model is put in Appendix C.
Table 11: Model and Training configuration. We report the number of layers (#Layer), hidden states dimension (Width), training tokens (Tokens)), global batch size in number of tokens (Batch), peaking learning rate (LR), and total training steps (Steps).
Name Size Width Layers Tokens Batch LR Steps Training FLOPs Transformer-s 110M 768 12 2.2B 0.5M 6.0e-4 4.2K 1.69e+18 Transformer-m 335M 1024 24 6.7B 0.5M 3.0e-4 13K 1.55e+19 Transformer-l 729M 1536 24 14.6B 0.5M 2.5e-4 28K 7.03e+19 Transformer-xl 1274M 2048 24 25.5B 0.5M 2.0e-4 49K 2.10e+20
@@
Efficiency To boost efficiency in training and inference, our code is based on PyTorch but incorporates fast CUDA kernels. We utilize Flash Attention [24], fast LayerNorm, and rotary embedding from TransformerEngine [53], and fuse operations including bias and GeLU. For our inference speed test, we use bfloat16. These techniques are applied to all models to ensure a fair latency or throughput comparison.
Wide and structured network To reduce the parameters and FLOPs of both attention and FFN, we adopt a straightforward approach. First, the dimension of the Key or Value in attention is set to 256 following [7]. Second, the rank and attention dimension are selected to be half of the hidden state dimension. This creates a structured network that looks like a bottleneck. We put the detailed configuration in Table 12.
Table 12: The intermediate dim. refers to the intermediate dimension of FFN. Attention dim. refers to the dimension of scaled-dot product. KV dim indicates the dimension used for KVCache, which is picked according to Team et al. [7].
Method hidden dim. intermediate dim. attention dim. KV dim. Transformer-m 1024 4096 1024 256 Transformer-m (GQA) 1024 4864 1024 256 Ours (R=512) 1024 4864 512 256 Transformer-l 1536 6144 1536 256 Transformer-l (GQA) 1536 7424 1536 256 Ours (R=768) 1536 7424 768 256

### v2

@@ B.4 Results on Refinedweb Dataset
Architecture Model FFN Model Training Loss PPL Size (M) Size (M) FLOPs (G) Tokens (B) FLOPs Transformer-s 110.0 56.6 262.9 2.2 1.69e+18 3.2569 25.97 LowRank 74.0 20.9 189.8 2.2 1.22e+18 3.3748 29.22 LowRank 2.2 1.39e+18 3.3329 28.02 LowRank 3.0 1.69e+18 3.2928 26.92 LowRank 2.2 1.69e+18 3.2866 26.75 BlockDense 74.0 20.9 189.8 2.2 1.22e+18 3.3731 29.17 BlockDense 2.2 1.39e+18 3.3338 28.04 BlockDense 3.0 1.69e+18 3.2982 27.06 BlockDense 2.2 1.69e+18 3.2856 26.73 BlockShuffle 74.0 20.9 189.8 2.2 1.22e+18 3.3994 29.95 BlockShuffle 2.2 1.39e+18 3.3583 28.74 BlockShuffle 3.0 1.69e+18 3.3218 27.71 BlockShuffle 2.2 1.69e+18 3.3011 27.14 Transformer-m 335.1 201.3 788.7 6.7 1.55e+19 2.9062 18.29 LowRank 202.4 68.7 517.0 6.7 1.01e+19 3.0251 20.60 LowRank 6.7 1.21e+19 2.9907 19.90 LowRank 10.2 1.54e+19 2.9359 18.84 LowRank 6.7 1.55e+19 2.9310 18.75 BlockDense 198.7 64.9 509.3 6.7 1.00e+19 3.0371 20.85 BlockDense 6.7 1.19e+19 3.0008 20.10 BlockDense 10.4 1.55e+19 2.9491 19.09 BlockDense 6.7 1.55e+19 2.9420 18.95 BlockShuffle 202.4 68.7 517.0 6.7 1.01e+19 3.0501 21.12 BlockShuffle 6.7 1.21e+19 3.0135 20.36 BlockShuffle 10.2 1.54e+19 2.9627 19.35 BlockShuffle 6.7 1.55e+19 2.9525 19.15 Transformer-l 729.1 453.0 1646.9 14.6 7.03e+19 2.6594 14.29 LowRank 430.7 154.5 1035.6 14.6 4.42e+19 2.7527 15.69 LowRank 23.3 7.03e+19 2.6917 14.76 LowRank 14.6 7.01e+19 2.6850 14.66 BlockDense 430.7 154.5 1035.6 14.6 4.42e+19 2.7570 15.75 BlockDense 23.3 7.03e+19 2.6946 14.80 BlockDense 14.6 7.01e+19 2.6941 14.79 BlockShuffle 430.7 154.5 1035.6 14.6 4.42e+19 2.7735 16.01 BlockShuffle 23.3 7.03e+19 2.7053 14.96 BlockShuffle 14.6 7.01e+19 2.7104 15.04 Transformer-xl 1274.1 805.3 2814.3 25.5 2.10e+20 2.5226 12.46 LowRank 743.6 274.7 1727.7 25.5 1.29e+20 2.6062 13.55 LowRank 41.5 2.10e+20 2.5464 12.76 LowRank 25.5 2.10e+20 2.5539 12.86 BlockDense 728.5 259.7 1696.8 25.5 1.27e+20 2.6204 13.74 BlockDense 42.2 2.10e+20 2.5590 12.92 BlockDense 25.5 2.10e+20 2.5637 12.98 BlockShuffle 743.6 274.7 1727.7 25.5 1.29e+20 2.6254 13.81 BlockShuffle 41.5 2.10e+20 2.5623 12.97 BlockShuffle 25.5 2.10e+20 2.5678 13.03
@@
Figure 14: Performance between dense Transformer, Structured FFN (BlockDense and BlockShuffle, LowRank in Fig. 8) with 32% parameters either trained with self-guided training or more tokens across four sizes. Circle size indicates model FLOPs. To enable straightforward comparison, we controlled their training FLOPs to be the same.
@@ Appendix C Implementation details
Model
Architecture details are provided in Table 10. We consider four baseline Transformer sizes, ranging from 110M to 1.3B parameters, with widths from 768 to 2048. For structured models, we first adopt two configurations as described in the main paper, reducing the FFN module to 63% and 32% of its original parameters by adjusting the rank $R$ and number of blocks $B$. Only $R$ and $B$ values specifically associated with structured matrices are modified, as seen in Table 6. Then, for more comparable results, we consider wide and structured networks, where the attention module is also structured by reducing the attention heads. We also present the transformer with GQA version [2] here, configured with 256 dimensions for the KVCache [9] and an enlarged FFN intermediate dimension following [8]. Based on this GQA version, we apply LowRank matrices to the FFN module with a rank half of the model or FFN width and use a smaller attention inner dimension to further reduce the parameters of the attention module. This allows us to maintain the parameter ratio between the attention and FFN modules.
Note that in all experiments, we do not apply structured matrices to the first FFN module, as doing so can lead to non-negligible performance loss in models on shallow networks. For dense models, we use Gaussian random weight initialization with a standard deviation of 0.02. For structured matrices, spectral initialization is applied for LowRank, and orthonormal initialization for the other two, based on initial experiments.
Table 10: Detailed configurations of the baseline Transformers, along with those using GQA [2] and wide and structured networks. The latter two are employed in the scaling study in Sec. 4.2. For other structured models which have 63% and 32% of the original FFN parameters, we adjust only the rank and number of blocks for each method and put the configuration directly in Table 6. Width denotes the model width or the input and output dimensions of the attention and FFN modules. Intermediate dim. refers to the intermediate dimension of the FFN. Attention dim. specifies the dimension used in scaled-dot product attention. KV dim. represents the dimension used for KVCache, as selected according to Team et al. [9].
Model Size (M) Layers Width Intermediate dim. Attention dim. KV dim. Transformer-s 110.0 12 768 3072 768 768 Transformer-s (GQA) 110.0 12 768 3584 768 256 Wide and Structured (R=384) 81.1 12 768 3584 512 256 Transformer-m 335.1 24 1024 4096 1024 1024 Transformer-m (GQA) 335.1 24 1024 4864 1024 256 Wide and Structured (R=512) 219.4 24 1024 4864 512 256 Transformer-l 729.1 24 1536 6144 1536 1536 Transformer-l (GQA) 729.1 24 1536 7424 1536 256 Wide and Structured (R=768) 464.4 24 1536 7424 768 256 Transformer-xl 1274.1 24 2048 8192 2048 2048 Transformer-xl (GQA) 1274.1 24 2048 9984 2048 256 Wide and Structured (R=1024) 799.6 24 2048 9984 1024 256
Dataset
@@
We use the RefinedWeb dataset [48], a carefully curated subset of CommonCrawl, optimized for filtering and deduplication, providing 600B tokens for public use. Due to its large size, we shuffle, extract, and tokenize it in advance. To manage memory efficiently, token IDs are stored using np.memmap, preventing the need to load all data into CPU memory simultaneously. The maximum sequence length is set to 1024. Following scaling laws [1], we allocate tokens at 20 times the number of parameters for each baseline model in all experiments, except for the 300B token training.
Table 11: Basic training configuration used in all experiments except for the overtraining regime with 300B tokens. Note that we apply the same global batch size (Batch) and the same peak learning rates (LR) to both dense and structured models to avoid hyperparameter search. The hyperparameter values are selected based on Zhang et al. [50], Gu and Dao [49].
Model Tokens Batch LR -s size Dense 2.2B 512 6.0e-4 Structured -m size Dense 6.7B 512 3.0e-4 Structured -l size Dense 14.6B 512 2.5e-4 Structured -xl size Dense 25.5B 512 2.0e-4 Structured
Table 12: Training configuration for 300B token training. Different studies [49, 52, 47, 51] employ very different learning rates in this setting, which also differ from training-compute scaling studies [1]. To avoid extensive tuning, we follow the hyperparameter scaling rule of Transformer proposed by Bi et al. [53], determining batch size and learning rate based on training FLOPs.
Model Size (M) Batch LR -s size Dense 110.0 1280 9.1e-4 Structured 81.1 1024 9.6e-4 -m size Dense 335.1 1792 7.8e-4 Structured 219.4 1536 8.4e-4 -l size Dense 729.1 2304 7.1e-4 Structured 464.4 2048 7.6e-4 -xl size Dense 1274.1 2816 6.6e-4 Structured 799.6 2304 7.1e-4
Training
We use A100 80G GPUs for training and evaluation, employing mixed precision (bfloat16 and float32) with torch.cuda.amp to accelerate training. Training FLOPs are calculated following Megatron [54], including all matrix multiplications.
Different hyperparameters are used for 300B token training and other experiments. For basic training in training FLOPs scaling and self-guided training studies, configurations are listed in Appendix C. Hyperparameter values are selected based on the scaling law studies of Zhang et al. [50], Gu and Dao [49], where we use the same learning rates and global batch size for both dense and structured models. Additional details include the AdamW optimizer with 0.1 weight decay, betas of [0.9, 0.999], cosine annealing learning rate scheduler with 10% linear warm-up, and 0.1$\times$ minimum value. Dropout is set to 0.0, and gradient clipping to 1.0.
In the overtraining regime where the training duration is super long, however, smaller betas [0.9, 0.98] are required for stable training, even for baseline Transformers. Previous studies [49, 52, 47, 51] adopt very different learning rates in this setting, differing from training-compute scaling studies [1]. To avoid extensive searching, we follow the hyperparameter scaling rule of Transformer proposed by Bi et al. [53], determining the global batch size and learning rate based on training FLOPs. Specifically, batch size is defined by $0.3118\times(\texttt{training FLOPs}^{-0.125})$, and learning rate by $0.2920\times(\texttt{training FLOPs}^{0.3271})$, giving the results in Appendix C. It can be seen that our wide and structured models trained on 300B tokens will use a slightly higher learning rate and smaller batch size compared to the larger Transformer.
Efficiency
@@
To enhance training and inference efficiency, our code is based on PyTorch but incorporates optimized CUDA kernels. We leverage Flash Attention [25], fast LayerNorm, and rotary embeddings from TransformerEngine [56], along with fused operations including bias and GeLU. For inference speed testing, we use bfloat16. These techniques are consistently applied to all models to ensure fair latency and throughput comparisons.
@@ Appendix D Broader Impacts
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: We describe our claims and contributions in Section 1.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We discuss limitations in Section 5.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: The paper does not include theoretical results.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We describe our experimental setup and implementation details in Section 4 Appendix C.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general, releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We open-source our code. The dataset we use is publicly available.
  25. Guidelines:
    • The answer NA means that the paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We describe our experimental setup in Section 4 and give details in Appendix C.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: Training large language models requires significant time and compute, making it infeasible to run multiple seeds. Nonetheless, results from single training runs are highly reliable and transferrable, as demonstrated by neural scaling laws. Presenting results without error bars in LLM research is standard practice.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: Computing resources are discussed throughout the paper.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: The research conducted in the paper conforms with the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: Societal impacts are discussed in Appendix D.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: The paper does not release data or models. not pose such risks.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: All the assets used have been cited.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: We provide documented code.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.